OpenStudy (anonymous):

The 5th term in a geometric sequence is 160. The 7th term is 40. What are the possible values of the 6th term of the sequence? (1 point) ±70 70 ±80 80

hartnn (hartnn):

Geometric sequence has the property that any 3 consecutive terms are in Geometric progression. so you can directly use the formula that \(\Large a_6^2 = a_5\times a_7\) just plug in and find 2 possible values of 6th term, ie., a6

hero (hero):

Perhaps it would be a wee bit clearer to write the notation as \[a_6 = \sqrt{(a_5 \times a_7)}\]